OpenStudy (anonymous):

Simplify (2-3i)(7-4i)

OpenStudy (lgbasallote):

use FOIL first then tell me what you got

OpenStudy (anonymous):

14-8i-21i+12i?

OpenStudy (lgbasallote):

you were right in the first three terms...last term is 12i^2 because it's i times i got it?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now what?

OpenStudy (lgbasallote):

substitute i^2 into -1

OpenStudy (anonymous):

now simplify since you know i^2 = -1

OpenStudy (anonymous):

2-29i?

OpenStudy (lgbasallote):

seems abt right

OpenStudy (anonymous):

\[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\] is another way to do it

OpenStudy (anonymous):

and the way you should really learn at some point

OpenStudy (anonymous):

Thank you for your help.

OpenStudy (lgbasallote):

@satellite73 loves memorizing formulas i see...i believe in the KISS principle so i just do it traditionally :P

OpenStudy (anonymous):

actually that is the definition not a formula if you see this topic in a later class that is what you will do, take my word for it

OpenStudy (lgbasallote):

but if you do forget it...FOIL/distributive property is a good thing to back up to